OpenStudy (anonymous):

The data represent the results for a test for a certain disease. Assume one individual from the group is randomly selected. Find the probability of getting someone who tests positive, given that he or she had the disease. The individual actually had the disease. Yes No Positive 141 6 Negative 11 142 @jim_thompson5910 I thought that I was done. Look this one. lol!! I am not sure what I am doing wrong!! Please help me!! Thanks!

jimthompson5910 (jim_thompson5910):

key phrase: given that he or she had the disease

jimthompson5910 (jim_thompson5910):

so we know for a fact that this person has the disease

jimthompson5910 (jim_thompson5910):

how many people total have the disease?

OpenStudy (anonymous):

It's not suppose to be 141+6= 147. Like the one before?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

that's the total who tested positive

jimthompson5910 (jim_thompson5910):

how many actually have the disease?

OpenStudy (anonymous):

ooh, 300

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

got it 152

jimthompson5910 (jim_thompson5910):

the first column has a "yes" under "The individual ACTUALLY had the disease"

jimthompson5910 (jim_thompson5910):

so the first column represent everyone who actually has the disease

OpenStudy (anonymous):

I got it... Oh boyy!!

OpenStudy (anonymous):

I am getting it now.. I was confused reading this.!! I got the answer.

OpenStudy (anonymous):

P(141/152)=0.9276=0.928 =D

jimthompson5910 (jim_thompson5910):

that's correct, but the notation is a bit off P(test positive | has disease) = (# of people who test positive and have disease)/(# of people who have disease) P(test positive | has disease) = 141/152

jimthompson5910 (jim_thompson5910):

The vertical bar | means "given"

OpenStudy (anonymous):

I understand now.. Yes I did it that way to round it!! But I am clear now... Thanks again for your help!! :*

jimthompson5910 (jim_thompson5910):

np